OpenStudy (anonymous):

lim as x -> 0+ for x/(ln(1-x^3))

OpenStudy (anonymous):

Do I use L'Hopital or should I do it normally?

OpenStudy (anonymous):

l'Hopital's since it it 0/0

OpenStudy (anonymous):

That gives me 0, agreed? Or am I missing something?

OpenStudy (anonymous):

0/0 does not equal zero. nor does the limit as x->0+ after applying lHopital's

OpenStudy (anonymous):

\[\frac{ f'(x) }{ g'(x) } = \frac{x^3 - 1 }{ 3x^2 }\]

OpenStudy (anonymous):

What do I do when I get to -1/0 then? I guess that's where I got confused and for some reason just wrote it as equal to 0..

OpenStudy (anonymous):

-1/0+ = - infinity = -1/0- you can check on your calculator that 1/±0.0000000001 = a big number. as the denominator gets infinitely smaller, 1/it approaches infinity

OpenStudy (anonymous):

Ahhhhh. Apply it vice subbing in 0 all the time, aka think not plug-n-chug. Thank you

OpenStudy (anonymous):

glad i could help :)